OpenStudy (anonymous):

i need help finding the first derivative of this problem Ln(3x^4+5)

OpenStudy (jamesj):

You need to use the chain rule here. Let g(x) = ln( 3x^4 + 5 ) = ln( f(x) ) where f(x) = 3x^4 + 5 Then...

OpenStudy (anonymous):

hold on before you continue that should have been 5x not just 5 sorry

OpenStudy (jamesj):

\[ \frac{dg}{dx} = \frac{dg}{df}\frac{df}{dx} \] See what to do now?

OpenStudy (jamesj):

Ok f(x) = 3x^4 + 5x

OpenStudy (anonymous):

right thats what I've been using and i keep coming up this (12x^3+5)/(3x^4+5) which is not the correct answer..

OpenStudy (jamesj):

No but, (12x^3+5)/(3x^4+5x) is

OpenStudy (jamesj):

...putting the 5x in the denominator, vs. 5

OpenStudy (anonymous):

OMG me forgetting that x is a serious problem lol thanks a ton!